A new pizzeria has opened near you home, the owner has decided to offer a 20% discount on

the opening day on the pizza whose original price is x dollars. The owner is offering additional

15% discount coupon to those who agree to fill a feedback form afterwards. You and your

friends have agreed to provide feedback and thus have become entitled to the additional 15%

discount coupon. Use the concept of composite functions to identify the amount that you will

pay after applying the coupon on the sale price.


1
Expert's answer
2021-09-27T07:31:20-0400

The price will be "(1-0.15)\times(1-0.2)\times x = 0.85\times\ 0.8\times\ x =0.68 x"

The discount, thus, is "1-0.68=0.32 = 32"%

Answer: The price is 0.68x dollars.